Session: May 2002 (session 2)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Xinyong Huang at Tue May 21 12:00:53 CEST 2002) Success

Medicina (orfei at Tue May 21 12:10:26 CEST 2002) Success

Torun (Kaz Borkowski at Tue May 21 12:12:29 CEST 2002) Minor failures
The sensitivity in the LCP channel was low (Tsys about 200 K) for yet unknown reason. No other known problems.

Jodrell_Bank (Alastair Gunn at Tue May 21 12:24:14 CEST 2002) Success
No known problems.

Cambridge (Alastair Gunn at Tue May 21 12:24:57 CEST 2002) Minor failures
Data only in BBCs 2 and 3; 1 and 4 failed. Investigating. No other known problems.

Effelsberg (A. Kraus at Tue May 21 13:44:24 CEST 2002) Minor failures
We could not start earlier than UT0924. No other known problems.

Urumqi (Liu Xiang at Wed May 22 03:59:37 CEST 2002) Success

Noto (C.Trigilio at Thu May 23 14:33:53 CEST 2002) Success

Westerbork (Tony F\;oley at Mon May 27 12:13:26 CEST 2002) Success
No known problems. All telescopes in array SEFD about 70Jy